OpenStudy (anonymous):

(-3a^3b^4)^3(a^3b^4)^3 What is the simplified form? Please Help Me, Thank You =)

OpenStudy (anonymous):

watchmath you must spend even more time here than i do.

OpenStudy (watchmath):

Cube each expression \((-3)^3\cdot a^{9}\cdot b^{12}\cdot a^9\cdot b^{12}=-27a^{18}b^{24}\)

OpenStudy (anonymous):

Thank You!!!!!!!!!! =)))))